# Large deviations in non-Markovian stochastic epidemics

### v1

@@ Abstract
III C. SIS model under non-Markovian infection
III.1 C1. Non-Markovian infection and Markovian recovery
@@
In this section we consider the complementary case of non-Markovian infection with Markovian or non-Markovian recovery. Let us start with Markovian (exponential) recovery and gamma-distributed infection. Here, the WTs satisfy
$\psi_{1}(t)=\frac{(\alpha\lambda_{1})^{\alpha}t^{\alpha-1}e^{-\alpha\lambda_{1}t}}{\Gamma(\alpha)},\quad\psi_{2}(t)=\lambda_{2}e^{-\lambda_{2}t},$ (S17)
with means of $\lambda_{1}=NR_{0}x_{s}x_{i}$ and $\lambda_{2}=Nx_{i}$.
We can now repeat the calculations done in the main text for this case and compute the quasistationary (QSD) distribution in the long-lived metastable state within the SIS model. In Fig. S1 we show how varying the shape parameter $\alpha$ affects the QSD’s mean and its standard deviation for gamma-distributed infection and exponential recovery. Normalization of the mean and standard deviation is performed as in Fig. 4 in the main text. Unlike the recovery case in Fig. 4, here increasing $\alpha$ leads to a larger standard deviation and a smaller mean. The disagreement with the theory here again occurs as the mean decreases towards 0 and the WKB theory becomes invalid.
@@
Figure S1: The normalized mean $x_{i}^{*}$ (a) and normalized standard deviation $\sigma$ (b) as functions of $\alpha$, for gamma-distributed infection and exponentially-distributed recovery, $N=1000$, $R_{0}=1.5$, and $10^{3}$ runs per $\alpha$. Simulations (symbols) are compared with theory (blue solid line). In all panels the initial conditions was $x_{i}(0)=x_{i}^{*}$.
III.2 C2. Non-Markovian infection and recovery
@@
One can also consider the case of non-Markovian infection and recovery, with WTs both distributed according to the gamma distribution. In this case the WTs satisfy
@@
Figure S2: The normalized mean $x_{i}^{*}$ (a) and normalized standard deviation $\sigma$ (b) as functions of $\alpha$, for gamma-distributed infection and gamma-distributed recovery, $N=5000$, $R_{0}=1.5$, and $10^{4}$ runs per $\alpha$. Simulations (symbols) are compared with theory (blue solid line). In all panels the initial condition was $x_{i}(0)=x_{i}^{*}$.
In Fig. S2 we show the dependence of the mean and standard deviation of the QSD on $\alpha$ (when both quantities are normalized as in Fig. 4 in the main text), for gamma-distributed infection and recovery with an identical shape parameter. Naturally, this choice is arbitrary; however, while a lengthy analysis can be performed by doing many calculations in the phase space of $(\alpha_{1},\alpha_{2})$, we merely wanted to show that an analysis where all the reactions are non-Markovian is feasible. Notably, even a simplified choice of having an identical $\alpha$ value, highlights some of the interesting effects that emerge. As expected, when both processes vary together, the mean remains constant; yet, the standard deviation changes qualitatively in a manner similar to the recovery case (Fig. 4 in the main text). This shows that, although the effects on the mean are exactly inverse for infection and recovery, their impact on fluctuations is not, as is evident from Eq. (12) in the main text.
IV Power-law waiting time distribution
In this section, we show that the asymptotic memory kernels can be obtained in a similar manner also for different WT distributions. Here we assume that infection is Markovian and recovery is non-Markovian, with a WT following a power-law distribution:
with means of $\lambda_{1}=NR_{0}x_{s}x_{i}$ and $\lambda_{2}=Nx_{i}$. Here, the shape parameter is bounded, $\alpha>1$, to ensure a finite mean, and in the limit $\alpha\to\infty$ the distribution converges to an exponential one. Notably, the regime of $\alpha>1$ in the gamma distribution does not exist here; namely, by choosing a power-law WT we can only increase the WT width and the corresponding memory strength.
Performing the explicit calculation of the normalized asymptotic memory kernel using section A, we find
where $E_{m}(z)\equiv\int_{1}^{\infty}e^{-z\tau}\tau^{-m}d\tau\,$ is the exponential integral function. At this point, one can plug these memory kernels into the late-time master equations and perform calculations along the same lines as in the main text, which allows to find the quantities of interest, e.g., outbreak-size distribution within the SIR model, or MTE within the SIS model.

### v2

@@ Abstract
We now provide a more detailed explanation of how to compute the standard deviation. As noted in the main text, we use the Gaussian approximation $\sigma\simeq\left|N\left.\mathcal{S}^{\prime\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}\right|^{-1/2}$. Compared to the SIS model, deriving $\mathcal{S}(x_{s}^{*})$ at $x_{s}^{*}=\bar{x}_{s}^{*}$ in the SIR model is more challenging for two reasons: the expressions for $\mathcal{S}(x_{s}^{*})$ and $\bar{x}_{s}^{*}$ are more complex, and there is an implicit function $m(x_{s}^{*})$ that must be accounted for in our derivation. To calculate $\left.\mathcal{S}^{\prime\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}$ we first need to find $\left.m^{\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}$ and $\left.m^{\prime\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}$. Rewriting Eq. (S15) as $F(m,x_{s}^{*})=0$, the derivatives of $m$ follow from the chain rule, $dF/dx_{s}^{*}=\left(\partial F/\partial m\right)\left(dm/dx_{s}^{*}\right)+\partial F/\partial x_{s}^{*}$ with $dF/dx_{s}^{*}=0$
$\left.m^{\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}$ $=$ $\!\!\left.-\frac{F_{x_{s}^{*}}}{F_{m}}\right|_{\bar{x}_{s}^{*}},$ (S18)
$\left.m^{\prime\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}$ $=$ $\!\!\left.-\frac{F_{m}^{2}F_{{x_{s}^{*}}{x_{s}^{*}}}\!-\!2F_{m}F_{x_{s}^{*}}F_{{x_{s}^{*}}m}\!+\!F_{x_{s}^{*}}^{2}F_{mm}}{F_{m}^{3}}\!\right|_{x^{*}_{s}=\bar{x}_{s}^{*}}\!\!\!\!,$
where $F_{z}=\partial F/\partial z$. Using these and the implicit formula for $\bar{x}_{s}^{*}$ (found by plugging $m(\bar{x}_{s}^{*})=1,x_{s}^{*}=\bar{x}_{s}^{*}$ in Eq. (S15)) we obtain a compact expression for $\left.\mathcal{S}^{\prime\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}$
$\left.\mathcal{S}^{\prime\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}=\frac{\left[\xi(\bar{x}_{s}^{*})-1\right]^{2}}{1-\bar{x}_{s}^{*}+I(\bar{x}_{s}^{*})},$ (S19)
where $\xi(\bar{x}_{s}^{*})=\mathcal{M}_{2}(\bar{x}_{s}^{*})/\mathcal{M}_{1}(\bar{x}_{s}^{*})$, $I(\bar{x}_{s}^{*})=\int_{\bar{x}_{s}^{*}}^{1}\xi(z)^{2}dz$. For $\alpha=1$, this reduces to $\xi(\bar{x}_{s}^{*})=1/(R_{0}\bar{x}_{s}^{*})$, $I(\bar{x}_{s}^{*})=-\left(1-\bar{x}_{s}^{*}\right)/\left(R_{0}^{2}\bar{x}_{s}^{*}\right)$, recovering the Markovian result [28]
$\left.\mathcal{S}^{\prime\prime}(x_{s}^{*})\right|_{\bar{x}_{s}^{*}}=\frac{\left(R_{0}\bar{x}_{s}^{*}-1\right)^{2}}{\left(1-\bar{x}_{s}^{*}\right)\bar{x}_{s}^{*}\left(R_{0}^{2}\bar{x}_{s}^{*}+1\right)}.$ (S20)
IV D. Quasi-stationary distribution in the non-Markovian SIS model
Figure S1: QSDs at $\alpha\!=\!0.5,1,2$ in (a)–(c) in the SIS model of epidemics, for gamma-distributed infection and Markovian recovery, with N=1000N\!=\!1000, $R_{0}\!=\!2$, and $10^{5}$ runs per $\alpha$. Here we compare simulations (symbols) to theory (solid line) and to a Markovian theory adjusted to the same mean (dashed line), with $x_{i}(0)=x_{i}^{*}$. The distributions are computed by averaging over a single trajectory from $t=50$ until $t=10^{5}$ (in units of $\gamma^{-1}$.
Here we compare our analytical prediction for the quasi-stationary distribution (QSD) [Eq. (13) in the main text] with numerical simulations, for the case of Markovian infection and non-Markovian recovery with gamma-distributed waiting times. In Fig. S1 we show excellent agreement between the simulated QSDs for various values of $\alpha$ and our theoretical prediction, in the case of non-Markovian, gamma-distributed infection, and Markovian recovery. We also plot by dashed lines the QSDs that would be obtained if we use a Markovian theory with adjusted reaction rates, obtained by replacing the gamma-distributed infection process with an exponential one, and adjusting the infection rate by $R_{0}\to R_{0}\alpha(2^{1/\alpha}-1)$, see main text. While being a good approximation close to the mean, for $\alpha\neq 1$ the tails are missed by this adjusted distribution as can be seen in panels (a) and (c). Notably, this effect strengthens as $\alpha$ is further decreased (or increased). This demonstrates that the non-Markovian effects cannot be fully captured by simply adjusting the mean of a Markovian reaction.
V E. SIS model under non-Markovian recovery
V.1 E1. SIS model under Markovian infection and Non-Markovian recovery
@@
This section is dedicated to the exploration of the SIS model in the complementary case of non-Markovian, gamma-distributed recovery with Markovian infection. The WTs satisfy
$\psi_{1}(t)=\lambda_{1}e^{-\lambda_{1}t},\quad\psi_{2}(t)=\frac{(\alpha\lambda_{2})^{\alpha}t^{\alpha-1}e^{-\alpha\lambda_{2}t}}{\Gamma(\alpha)},$ (S21)
with means of $\lambda_{1}=NR_{0}x_{s}x_{i}$ and $\lambda_{2}=Nx_{i}$. Performing the explicit calculation of the normalized asymptotic memory kernels as in Sec. A, we find
$\mathcal{M}_{1}=R_{0}x_{i}x_{s},\quad\mathcal{M}_{2}=\frac{R_{0}x_{i}x_{s}}{\left(1+R_{0}x_{s}/\alpha\right)^{\alpha}-1}$ (S22)
Using these memory kernels, we can compute the mean and standard deviation of the QSD, within the SIS model, under Markovian infection and non-Markovian recovery. In Fig. S2(a,b) we show how the shape parameter $\alpha$ affects the metastable mean and its standard deviation, both normalized by their Markovian values ($x_{i}^{*}(1)=1/3$, $\sigma(1)\approx 0.026$ for $R_{0}=1.5$). As $\alpha$ decreases, the mean drastically decreases and the variance increases; this is since the typical time between recovery events gets shorter as the WT distribution becomes wider and more skewed. The opposite occurs for $\alpha>1$; for $\alpha\to\infty$, the mean approaches $x_{i}^{*}(\infty)/x_{i}^{*}(1)\simeq(1-\ln 2/R_{0})/(1/3)\simeq 1.614$. Near the bifurcation, the mean vanishes and the WKB approximation breaks down.
The dependence of the variance on $\alpha$ as observed in Fig. S2(b) can be explained by looking at the right-hand-side of Eq. (14) in the main text, where the standard deviation $\sigma$ is expressed as a function of the mean $x_{i}^{*}$ and $\alpha$ (instead of $R_{0}$ and $\alpha$). Similarly as in the case of non-Markovian infection and Markovian recovery, the dominating factor in shaping $\sigma$ is its dependence on $x_{i}^{*}$, $\sigma\sim\sqrt{1-x_{i}^{*}}$. However, $\sigma$ is also affected by non-Markovianity: $\sigma\sim\left[2\alpha(1-2^{-1/\alpha})\right]^{-1/2}$, which reinforces the dependence on $x_{i}^{*}$ in this case. Notably, the WKB approximation breaks down as $x_{i}^{*}$ approaches 0, as seen in the figure.
@@
Figure S2: The normalized mean $x_{i}^{*}$ (a) and normalized standard deviation $\sigma$ (b) as functions of $\alpha$, in the SIS model of epidemics, for exponential infection and gamma-distributed recovery, $N=5000$, $R_{0}=1.5$, and $10^{4}$ runs per $\alpha$. Simulations (symbols) are compared with theory (blue solid line). In all panels the initial condition was $x_{i}(0)=x_{i}^{*}$.
V.2 E2. SIS model under Non-Markovian infection and recovery
@@
Here, we consider the full non-Markovian case of having both processes of infection and recovery with gamma-distributed WTs. In this case the WTs satisfy
@@
In Fig. S3 we show the dependence of the mean and standard deviation of the QSD on $\alpha$ (where both quantities are normalized as in Fig. S2 and Fig. 3 in the main text), for gamma-distributed infection and recovery with an identical shape parameter. Naturally, this choice is arbitrary; however, while a lengthy analysis can be performed by doing many calculations in the phase space of $(\alpha_{1},\alpha_{2})$, we merely wanted to show that an analysis where all the reactions are non-Markovian is feasible. Notably, even a simplified choice of having an identical $\alpha$ value, highlights some of the interesting effects that emerge. As expected, when both processes vary together, the mean remains constant; yet, the standard deviation changes qualitatively in a manner similar to the recovery case (see Fig. S2). This shows that, although the effects on the mean are exactly inverse for infection and recovery, their impact on fluctuations is not, as is evident from Eq. (14) in the main text.
Figure S3: The normalized mean $x_{i}^{*}$ (a) and normalized standard deviation $\sigma$ (b) as functions of $\alpha$, in the SIS model of epidemics, for gamma-distributed infection and gamma-distributed recovery, $N=5000$, $R_{0}=1.5$, and $10^{4}$ runs per $\alpha$. Simulations (symbols) are compared with theory (blue solid line). In all panels the initial condition was $x_{i}(0)=x_{i}^{*}$.